# Language Models Surface the Unwritten Code of Science and Society

### v4

@@ Appendix H Details of the dataset of conference submissions
This dataset includes 26,731 submissions to 45 conferences: 1st ContinualAI Unconference, AAAI Conference on Artificial Intelligence 2024 and 2025, Symposium on Advances in Approximate Bayesian Inference 2024, Conference on Language Modeling 2024, Cooking Robotics Workshop 2024, Conference on Robot Learning 2023 and 2024, Conference on Parsimony and Learning 2024, IEEE/CVF Conference on Computer Vision and Pattern Recognition 2023 and 2024, Workshop on Distributed Infrastructure for Common Good 2023, Workshop on Embodiment-Aware Robot Learning 2024, European Conference on Computer Vision 2024, Conference on Empirical Methods in Natural Language Processing 2023, European Space Power Conference 2023, Fast, Low-resource, and Accurate Organ and Pan-cancer Segmentation in Abdomen CT 2023, SIGIR Workshop on Generative Information Retrieval 2024, ACM/IEEE International Conference on Human-Robot Interaction 2023 and 2024, International Conference on Integration of Science and Technology for Sustainable Development 2024, International Conference on Learning Representations 2017 to 2024, International Conference on Machine Learning 2023 and 2024, ACM International Conference on the Theory of Information Retrieval 2024, International Joint Conference on Artificial Intelligence 2024, International Semantic Web Conference 2024, ACM SIGKDD Conference on Knowledge Discovery and Data Mining 2023 and 2024, ACM International Conference on Multimedia 2024, Conference on Neural Information Processing Systems 2021-2024, Neuro-Symbolic Learning and Reasoning in the era of Large Language Models 2024, Next-generation Data Governance Workshop 2024, Tsinghua University Advanced Machine Learning 2024.
@@ Appendix I Prompts used in the paper
The dictionary method used in the paper to complement LLM labeling/matching is shown in Table 2. Note that this heuristic dictionary-like approach, while widely used in social science [20, 17], ultimately serves as a substitute rather than the major methods of labeling human comments of "mention" or matching priors and posteriors, presented in the paper.
@@
To address potential position bias, we adopt two mitigation strategies in the original paper. First, we construct paper pairs with substantial quality differences, as prior research has shown that position bias is most pronounced when the quality gap is small [36, 53]. Second, we randomize the positions of Paper 1 and Paper 2 across voting, while ensuring logically consistent prompts and vote aggregation. We then assess whether position bias is indeed stronger when the quality signal gap is small. Specifically, we regress the consistency rate — defined as the proportion of consistent judgments across 20 hypotheses after swapping paper positions — on the review score gap between the two papers. Our analysis (Table 4) shows that a larger review score gap is indeed associated with greater consistency, and the effect is statistically significant, though weak in magnitude.

### v5

@@ Appendix H Details of the dataset of conference submissions
This dataset includes 26,731 submissions to 46 conferences: 1st ContinualAI Unconference, AAAI Conference on Artificial Intelligence 2024 and 2025, Symposium on Advances in Approximate Bayesian Inference 2024, Conference on Language Modeling 2024, Cooking Robotics Workshop 2024, Conference on Robot Learning 2023 and 2024, Conference on Parsimony and Learning 2024, IEEE/CVF Conference on Computer Vision and Pattern Recognition 2023 and 2024, Workshop on Distributed Infrastructure for Common Good 2023, Workshop on Embodiment-Aware Robot Learning 2024, European Conference on Computer Vision 2024, Conference on Empirical Methods in Natural Language Processing 2023, European Space Power Conference 2023, Fast, Low-resource, and Accurate Organ and Pan-cancer Segmentation in Abdomen CT 2023, SIGIR Workshop on Generative Information Retrieval 2024, ACM/IEEE International Conference on Human-Robot Interaction 2023 and 2024, International Conference on Integration of Science and Technology for Sustainable Development 2024, International Conference on Learning Representations 2017 to 2025, International Conference on Machine Learning 2023 and 2024, ACM International Conference on the Theory of Information Retrieval 2024, International Joint Conference on Artificial Intelligence 2024, International Semantic Web Conference 2024, ACM SIGKDD Conference on Knowledge Discovery and Data Mining 2023 and 2024, ACM International Conference on Multimedia 2024, Conference on Neural Information Processing Systems 2021-2024, Neuro-Symbolic Learning and Reasoning in the era of Large Language Models 2024, Next-generation Data Governance Workshop 2024, Tsinghua University Advanced Machine Learning 2024.
@@ Appendix I Prompts used in the paper
The dictionary method used in the paper to complement LLM labeling/matching is shown in Table 2. Note that this heuristic dictionary-like approach, while widely used in social science (Guo et al., 2016; Frankel et al., 2022), ultimately serves as a substitute rather than the major methods of labeling human comments of "mention" or matching priors and posteriors, presented in the paper.
@@
To address potential position bias, we adopt two mitigation strategies in the original paper. First, we construct paper pairs with substantial quality differences, as prior research has shown that position bias is most pronounced when the quality gap is small (Li et al., 2024b; Shi et al., 2024). Second, we randomize the positions of Paper 1 and Paper 2 across voting, while ensuring logically consistent prompts and vote aggregation. We then assess whether position bias is indeed stronger when the quality signal gap is small. Specifically, we regress the consistency rate — defined as the proportion of consistent judgments across 20 hypotheses after swapping paper positions — on the review score gap between the two papers. Our analysis (Table 4) shows that a larger review score gap is indeed associated with greater consistency, and the effect is statistically significant, though weak in magnitude.